# KernelGPT: Enhanced Kernel Fuzzing via Large Language Models

### v2

@@ 3.2. Specification Validation and Repair
In this phase, inspired by recent work on LLM-based program repair (chatrepair, ), our goal is to validate the specifications generated by KernelGPT and automatically repair the invalid ones. This step is crucial, as LLMs may occasionally make mistakes during the description generation process. To address potential inaccuracies, we employ off-the-shelf validation tools. These tools analyze the specifications and provide error messages if discrepancies are found. Initially, KernelGPT uses the error messages from them to pinpoint inaccuracies in specific descriptions, effectively matching each error message to its corresponding description. Then, for those descriptions identified with errors, KernelGPT queries LLMs for correction, guided by few-shot examples. This process involves supplying LLMs with the incorrect description, the associated error messages, and relevant source code from the kernel codebase to repair. LLMs are then expected to output the correct descriptions.
@@ 4. Implementation
Specification generation. Rather than generating descriptions for all possible drivers and sockets, we focus on those not covered by existing specifications, which are often less thoroughly tested. We select drivers and sockets activated in our configuration, excluding ones used for debugging (e.g., /dev/gup_test) or requiring specific hardware/architecture, as testing them would be meaningless or impractical.
@@ 5.1.2. Coverage Improvement by New Specifications
To show the effectiveness of the new specifications synthesized by KernelGPT for kernel fuzzing, we integrate them with the existing Syzkaller specifications, resulting in a combined suite (Syzkaller + KernelGPT). We conduct a 24-hour fuzzing session (192 CPU hours). For comparison, we also run the original Syzkaller and a combined suite of Syzkaller with the descriptions generated by SyzDescribe (Syzkaller + SyzDescribe), each under identical conditions and with three repetitions. The results are depicted in Table 3, where the Syzkaller + KernelGPT suite covers 4,750 and 8,039 more lines of code than the original Syzkaller and the SyzDescribe integration, respectively. Additionally, KernelGPT adds 20,742 unique lines to Syzkaller’s coverage, in contrast to SyzDescribe’s contribution of 14,585 additional unique lines. These results underscore the contribution of KernelGPT-generated missing descriptions to enhancing coverage in kernel fuzzing.
@@ 5.2.1. Device Drivers
Table 5 presents the results for drivers. Due to space constraints, we omit the average number of unique crashes for each driver in Table 5. In total, KernelGPT triggers 24.0 unique crashes, compared to 21.0 by Syzkaller and 20.7 by SyzDescribe. Two baseline drivers, ashmem and fd#, are no longer supported in Linux 6 (“N/A”). Notably, KernelGPT achieves the highest line coverage and crashes, surpassing the baselines by at least 18.0% and 17.6%. Moreover, KernelGPT performs the best on 20 of 28 (excluding the 2 invalid drivers, highlighted in bold), whereas Syzkaller and SyzDescribe lead in only 4 and 4, respectively. This highlights the effectiveness of KernelGPT-generated specifications in enhancing fuzzing. For kvm driver (kvm, ), KernelGPT identifies two additional operation handlers, kvm_vm_fops and kvm_vcpu_fops, as dependencies. This leads to a coverage increase of 42.5% and 65.2% compared to baselines.
@@ 5.2.2. Sockets
Table 6 presents the results for sockets. We observe that KernelGPT can cover 18.6% more lines than our baseline Syzkaller, demonstrating the effectiveness of our generated specifications for sockets. Notably, KernelGPT describes significantly more syscalls than Syzkaller, showcasing KernelGPT’s capability in syscall discovery. This is also because Syzkaller often uses a single syscall with various command values, while KernelGPT generates distinct syscalls for each command value. For example, KernelGPT synthesizes 99 descriptions for the l2tp_ip6 socket handler, compared to Syzkaller’s 38. This is because the single Syzkaller syscall of this socket, getsockopt\$inet6_int, uses flags[inet6_option_types_int] as the command value list, encompassing 45 unique syscall identifier values.
@@ 6.1. Kernel Fuzzing
6.2. LLMs for Testing
Recent advancements have shown that LLMs (chatgpt, ; gpt4, ; starcoder, ; codex, ; codellama, ) excel in a variety of natural language processing (brown2020language, ) and programming tasks (xu2022systematic, ; bubeck2023sparks, ), including code comprehension and document generation. Their proficiency in diverse tasks is attributed to the extensive training on vast datasets, e.g., GPT4 (gpt4, ) is pre-trained using trillions of text tokens from the entire Internet. As a result, LLMs can be employed in various tasks simply by following instructions (ouyang2022training, ; bang2023multitask, ; chatgpt, ), eliminating the need for specialized training.
A growing body of research has focused on leveraging LLMs for testing, such as unit test generation (nie2023learning, ; lemieux2023codamosa, ; schafer2023adaptive, ; yuan2023no, ), fuzzing (fuzz4all, ; titanfuzz, ; deng2023large, ; yang2023white, ; hu2023augmenting, ; chatafl, ) and static analysis (li2023assisting, ). TitanFuzz (titanfuzz, ) is the pioneering work in applying modern LLMs for both generation-based (csmith, ; yarpgen, ) and mutation-based (donaldson2017automated, ; le2014compiler, ; wei2022free, ) fuzzing, specifically targeting machine learning systems. Fuzz4All (fuzz4all, ) further demonstrates that the multilingual potentials of LLMs can be utilized to serve as a universal fuzzer for a wide range of software systems. Existing works in utilizing LLMs in testing mainly focus on using them to create the test inputs for the systems under test directly, e.g., generating C programs for testing C compilers. By contrast, KernelGPT integrates the potential of LLMs with established fuzzing frameworks by synthesizing their components (i.e., input generators), complementary to existing LLM-based fuzzing techniques and offers additional benefits by leveraging the expertise and resources invested in well-developed fuzzing tools. Furthermore, to the best of our knowledge, KernelGPT is the first work to leverage LLMs for kernel fuzzing and demonstrate their ability to generate high-quality syscall specifications.

### v3

@@ Abstract.
Linux Kernel, Fuzzing, Large Language Models, Code Analysis
journalyear: 2025copyright: ccconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 30-April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’25), March 30-April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3676641.3716022isbn: 979-8-4007-1079-7/25/03ccs: Security and privacy Operating systems securityccs: Software and its engineering Software testing and debugging
@@ 3.2. Specification Validation and Repair
In this phase, inspired by recent work on LLM-based program repair (chatrepair, ), our goal is to validate the specifications generated by KernelGPT and automatically repair the invalid ones. This is because LLMs may occasionally make mistakes during the description generation process. To address potential inaccuracies, we employ off-the-shelf validation tools. These tools analyze the specifications and provide error messages if discrepancies are found. Initially, KernelGPT uses the error messages from them to pinpoint inaccuracies in specific descriptions, effectively matching each error message to its corresponding description. Then, for those descriptions identified with errors, KernelGPT queries LLMs for correction, guided by few-shot examples. This process involves supplying LLMs with the incorrect description, the associated error messages, and relevant source code from the kernel codebase to repair. LLMs are then expected to output the correct descriptions.
Existing validation tools are limited in their ability to detect semantic errors (syzlang, ), primarily focusing on syntax validation and simple semantic checks. For example, runtime validation of semantic correctness remains a significant challenge, which is why current syscall specification generation approaches do not incorporate it (syzdescribe, ; difuze, ). To ensure a more thorough evaluation, we manually examined the generated specifications (§ 5.1.3), demonstrating that KernelGPT successfully synthesizes semantically correct specifications.
@@ 4. Implementation
We set the default value of MAX_ITER to 5 (Line 1, Algorithm 1). For efficiency, our implementation caches and reuses results from previously explored paths. With these configurations, we observed no termination issues throughout our experiments.
Specification generation. Rather than generating descriptions for all possible drivers and sockets, we focus on those not covered by existing specifications, which are often less thoroughly tested. We select drivers and sockets activated in our configuration, excluding ones used for debugging (e.g., /dev/gup_test) or requiring specific hardware/architecture, as testing them would be meaningless or impractical. This filtering process is largely automated: debug drivers are easily identified by their _test suffix, and hardware-specific drivers can be filtered by focusing only on bootable modules.
@@ 5.1.1. Statistics of New Specifications
KernelGPT takes 4.7 hours to generate specifications for these 532 syscalls, more efficient than SyzDescribe, which requires 3.8 hours for just 146 syscalls. It processes approximately 5.56 million input tokens and generates 400,000 output tokens, with an average of 2,630 input and 189 output tokens per prompt. The total cost of  \$34 is negligible considering these specifications guide extensive fuzzing campaigns that typically run for days or weeks.
@@ 5.1.2. Coverage Improvement by New Specifications
To show the effectiveness of the new specifications synthesized by KernelGPT for kernel fuzzing, we integrate them with the existing Syzkaller specifications, resulting in a combined suite (Syzkaller + KernelGPT). We conduct a 24-hour fuzzing session (192 CPU hours). For comparison, we also run the original Syzkaller and a combined suite of Syzkaller with the descriptions generated by SyzDescribe (Syzkaller + SyzDescribe), each under identical conditions and with three repetitions. The results are depicted in Table 3, where the Syzkaller + KernelGPT suite covers 4,750 and 8,039 more basic blocks than the original Syzkaller and the SyzDescribe integration, respectively. Additionally, KernelGPT adds 20,742 unique basic blocks to Syzkaller’s coverage, in contrast to SyzDescribe’s contribution of 14,585 additional unique blocks. These results underscore the contribution of KernelGPT-generated missing descriptions to enhancing coverage in kernel fuzzing.
@@ 5.2.1. Device Drivers
Table 5 presents the results for drivers. Due to space constraints, we omit the average number of unique crashes for each driver in Table 5. In total, KernelGPT triggers 24.0 unique crashes, compared to 21.0 by Syzkaller and 20.7 by SyzDescribe. Two baseline drivers, ashmem and fd#, are no longer supported in Linux 6 (“N/A”). Notably, KernelGPT achieves the highest basic block coverage and crashes, surpassing the baselines by at least 18.0% and 17.6%. Moreover, KernelGPT performs the best on 20 of 28 (excluding the 2 invalid drivers, highlighted in bold), whereas Syzkaller and SyzDescribe lead in only 4 and 4, respectively. This highlights the effectiveness of KernelGPT-generated specifications in enhancing fuzzing. For kvm driver (kvm, ), KernelGPT identifies two additional operation handlers, kvm_vm_fops and kvm_vcpu_fops, as dependencies. This leads to a coverage increase of 42.5% and 65.2% compared to baselines.
@@ 5.2.2. Sockets
Table 6 presents the results for sockets. We observe that KernelGPT can cover 18.6% more basic blocks than our baseline Syzkaller, demonstrating the effectiveness of our generated specifications for sockets. Notably, KernelGPT describes significantly more syscalls than Syzkaller, showcasing KernelGPT’s capability in syscall discovery. This is also because Syzkaller often uses a single syscall with various command values, while KernelGPT generates distinct syscalls for each command value. For example, KernelGPT synthesizes 99 descriptions for the l2tp_ip6 socket handler, compared to Syzkaller’s 38. This is because the single Syzkaller syscall of this socket, getsockopt\$inet6_int, uses flags[inet6_option_types_int] as the command value list, encompassing 45 unique syscall identifier values.
@@ 6.1. Kernel Fuzzing
6.2. Learning-Based Fuzzing
Machine learning approaches to input generation (godefroid2017learn, ; liu2019deepfuzz, ; Rajpal2017NotAB, ; Sablotny2018RecurrentNN, ; cummins2018compiler, ) explored using sequence-to-sequence models to learn program syntax and generation patterns. For instance, Learn&Fuzz (godefroid2017learn, ) leveraged sequence-to-sequence models for grammar-based fuzzing by learning from sample inputs. DeepFuzz (liu2019deepfuzz, ) extended this approach to generate C programs for compiler testing. DeepSmith (cummins2018compiler, ) also demonstrated the potential of learning program patterns from a large corpus. However, these approaches faced significant limitations: they required extensive domain-specific training data, exhibited lower throughput compared to traditional fuzzers, and struggled to leverage existing fuzzing infrastructure. Our initial explorations show that even using more powerful LLMs for direct input generation in kernel fuzzing performed notably worse than basic Syzkaller.
Recent advancements have shown that LLMs (chatgpt, ; gpt4, ; starcoder, ; codex, ; codellama, ) excel in a variety of natural language processing (brown2020language, ) and programming tasks (xu2022systematic, ; bubeck2023sparks, ; xia2024agentless, ; xia2023automated, ). Their proficiency in diverse tasks is attributed to the extensive training on vast datasets, e.g., GPT4 (gpt4, ) is pre-trained using trillions of text tokens from the entire Internet. As a result, LLMs can be employed in various tasks simply by following instructions (ouyang2022training, ; bang2023multitask, ; chatgpt, ), eliminating the need for specialized training. Recently, a growing body of research has focused on leveraging LLMs for software testing, covering both unit test generation (nie2023learning, ; lemieux2023codamosa, ; schafer2023adaptive, ; yuan2023no, ) and fuzzing (fuzz4all, ; titanfuzz, ; yang2023white, ; hu2023augmenting, ; chatafl, ). For example, TitanFuzz (titanfuzz, ) pioneered the application of modern LLMs for both generation-based (csmith, ; yarpgen, ) and mutation-based (donaldson2017automated, ; le2014compiler, ) fuzzing, while Fuzz4All (fuzz4all, ) further demonstrated that the multilingual potentials of LLMs can be utilized to serve as a universal fuzzer for a wide range of software systems.
KernelGPT takes a fundamentally different approach from both traditional ML-based and recent LLM-based fuzzing techniques. Instead of directly generating test inputs, we integrate LLMs with mature fuzzing frameworks by synthesizing their components (i.e., input generators). This strategy leverages the expertise and resources invested in well-developed fuzzing tools while capitalizing on LLMs’ capabilities. To the best of our knowledge, KernelGPT is the first work to successfully apply LLMs to kernel fuzzing, demonstrating state-of-the-art performance in generating high-quality syscall specifications.
@@ 7. Conclusion
Acknowledgments
We are grateful to the anonymous reviewers and our shepherd, Youngjin Kwon, for their valuable feedback that helped improve this paper. This work was partially supported by NSF grant CCF-2131943 and Kwai Inc. Chenyuan Yang was partially supported by Boeing for research on Linux kernel testing. We also thank Ziqi Zhang for his helpful suggestions on the manuscript.